OpenStudy (ltrout):

[Fan + Medal] Find the real or imaginary solutions to the equations: 4x^2=-4x-1 2x^2+3=4x

OpenStudy (faiqraees):

apply the quadratic formula

OpenStudy (ltrout):

@FaiqRaees Could you provide an example using the first problem and I'll try it on the second?

OpenStudy (faiqraees):

General form of a quadratic \[\large\rm ax^2 + bx+c = 0\] \[\large\rm x = \frac {-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (faiqraees):

Can you apply it?

OpenStudy (ltrout):

@FaiqRaees So I should first change the problems to fit the ax^2+bx+c = 0 correct?

OpenStudy (foreveralone01):

^the ax^2+bx+c=0 is how you should set it up. Sorry if that didn't make any sense. So just insert your equation

OpenStudy (ltrout):

So I just leave the 4x^2=-4x-1 like it is?

OpenStudy (foreveralone01):

Yep

OpenStudy (ltrout):

But then there isn't a C on either side? This is confusing.

OpenStudy (foreveralone01):

True and hold on give me a sec

OpenStudy (faiqraees):

no arrange it

OpenStudy (faiqraees):

in that way you would know what will be the values of a b and c

OpenStudy (foreveralone01):

Okay just go with what they're saying because I'm not making sense lol